# spacing=nonfrench

### v1

@@ V-D Adaptive Graph-Guided Retrieval Performance
GPT-4 85.7±1.2% 87.0±0.9% 8.5±2.1% 12.3±1.8% 68±2.3%
GPT-4-Turbo 87.1±1.0% 88.2±0.8% 9.2±1.9% 14.1±1.6% 71±2.0%
Claude-3-Opus 84.9±1.3% 86.1±1.1% 7.8±2.3% 11.7±2.0% 67±2.4%
Claude-3-Sonnet 83.2±1.4% 85.2±1.2% 7.3±2.4% 10.9±2.1% 66±2.5%
Gemini-1.5-Pro 88.3±0.9% 89.1±0.7% 11.2±1.7% 15.8±1.5% 74±1.8%
CodeT5+ 86.5±1.1% 84.8±1.3% 10.6±1.8% 13.2±1.7% 72±1.9%
Chronos 90.2±0.6%*** 88.9±0.5%* 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
*p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001 compared to best baseline (two-tailed t-test)
TABLE V: Performance across code synthesis and debugging tasks (mean ± std over 5 runs). Statistical significance shown.
V-E Comparison with Agentic Code Tools
@@
Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
Cursor Auto IDE only None No Limited No 4.2%
Claude Code 200K tokens Session No Yes No 6.8%
OpenAI Codex 8K tokens None No No No 3.1%
Gemini CLI 1M tokens None No Yes Limited 9.7%
GitHub Copilot X 16K tokens None No Limited No 5.3%
Chronos Unlimited* Persistent Yes Yes Yes 65.3%
TABLE VI: Comparison of Chronos with agentic code tools. *Unlimited via intelligent retrieval and memory.
@@
V-F Cycle-Aware Debugging Case Study
@@
GPT-4ClaudeCodeT5+Chronos0022224444666688886.56.56.56.56.86.86.86.85.15.15.15.12.22.22.22.2ModelAvg. Cycles to Fix
@@
Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles—underscoring its benefit for continuous, automated codebase reliability.
V-G Case Studies
@@
V-G1 Case Study 1: Cross-Module Null Pointer Exception
@@
    Retrieved recent commits mentioning ”authentication” (found 3 relevant commits)
@@
    Traced data flow from auth module to export module
@@
    Identified missing null check after auth token refresh
@@
    Located similar pattern in 2 other modules
@@
Null Pointer Fix in Authentication Flow // In AuthService.java public AuthToken refreshToken(String userId) { AuthToken token = tokenCache.get(userId); if (token == null $|$ token.isExpired()) { token = authProvider.generateToken(userId); + if (token != null) { + tokenCache.put(userId, token); + } } return token; } // In ExportService.java public void exportUserData(String userId) { AuthToken token = authService.refreshToken(userId); + if (token == null) { + throw new AuthenticationException( + "Failed to authenticate user for export"); + } // ... rest of export logic }
Validation Results: All 47 test cases passed, including 3 new tests Chronos generated for the edge case.
V-G2 Case Study 2: Async Race Condition in Message Queue
@@
Chronos Analysis:
  • Retrieved message queue implementation and recent load test results
  • Identified race condition between acknowledgment and connection pooling
  • Found similar issue fixed 8 months ago in different component
  • Applied analogous solution with adaptations
@@
Race Condition Fix in Message Queue // MessageProcessor.java private void processMessage(Message msg) { try { handler.process(msg); - connectionPool.returnConnection(conn); - msg.acknowledge(); + // Fix: Acknowledge before returning connection + msg.acknowledge(); + connectionPool.returnConnection(conn); } catch (Exception e) { + // Ensure connection returned even on error + connectionPool.returnConnection(conn); msg.nack(); throw e; } }
@@
V-H Long-Context Debugging Performance
@@
GPT-4-32K 32K tokens 7.2±1.9% 3.1±1.2% 5.8±1.6% 5.4±1.5%
Claude-3-200K 200K tokens 9.8±2.1% 4.7±1.4% 8.3±1.8% 7.6±1.7%
Gemini-1.5-Pro-1M 1M tokens 14.3±2.3% 6.2±1.5% 11.7±2.0% 10.7±1.9%
Chronos Unlimited 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
***p ¡ 0.001 compared to Gemini-1.5-Pro-1M (paired t-test, n=100 tasks per category)
TABLE VII: Performance on debugging tasks requiring extensive context. Statistical significance demonstrates retrieval superiority.
The results demonstrate that raw context size alone cannot solve debugging. Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 6x.
V-I Detailed Performance Analysis
@@
GPT-4 82.3% 12.1% 3.2% 5.7% 18.9% 7.4%
Claude-3-Opus 79.8% 10.7% 2.8% 4.3% 16.2% 6.1%
Gemini-1.5-Pro 85.1% 15.3% 4.1% 6.9% 22.4% 9.8%
@@
TABLE VIII: Success rates by bug category. Chronos shows particular strength in complex bug types that require deep understanding.
Repository Size ¡10K LOC 10K-100K 100K-1M ¿1M LOC
GPT-4 15.2% 9.8% 4.3% 1.2%
Claude-3-200K 17.8% 11.2% 5.7% 2.1%
Gemini-1.5-Pro 21.3% 14.7% 8.9% 3.8%
@@
TABLE IX: Debugging success rates by repository size, demonstrating Chronos’s scalability.
V-J Multi-Code Association Retrieval Performance
@@
TABLE X: Multi-code association retrieval performance across different debugging contexts.
V-K Computational Efficiency
@@
Metric GPT-4 Claude-3 Gemini-1.5 Chronos Human Dev
Avg. Time to Fix 82.3s 76.9s 71.2s 134.7s 2.4 hours
@@
Cost per Bug \$0.47 \$0.52 \$0.68 \$0.89 \$180
Success Rate 8.5% 7.8% 11.2% 65.3% 94.2%
Effective Cost* \$5.53 \$6.67 \$6.07 \$1.36 \$191
@@
TABLE XI: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
V-K1 Inference Time Breakdown
@@
V-K2 Return on Investment Analysis
@@
VI DISCUSSION AND ABLATION ANALYSIS
VI-A Qualitative Insights

### v2

@@ V-G2 Oracle Retriever Experiments: Upper Bound Analysis
GPT-4o 87.8±1.0% 88.5±0.8% 10.3±1.9% 14.7±1.6% 72±2.0%
GPT-4.1 91.2±0.8% 90.7±0.7% 13.8±1.7% 18.2±1.5% 76±1.8%
Claude 4 Opus 92.8±0.7% 91.3±0.6% 14.2±1.6% 19.1±1.4% 78±1.7%
Claude 4 Sonnet 92.1±0.8% 90.9±0.7% 13.6±1.7% 18.5±1.5% 77±1.8%
DeepSeek V3 90.5±0.9% 89.8±0.8% 12.1±1.8% 16.3±1.6% 75±1.9%
Qwen2.5-32B 89.7±1.0% 88.9±0.9% 11.5±1.9% 15.7±1.7% 73±2.0%
Gemini 2.5 Pro 91.6±0.8% 90.2±0.7% 13.9±1.7% 17.9±1.5% 76±1.8%
Chronos 90.2±0.6%NS 88.9±0.5%NS 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
NS: Not significant, *p <0.05<0.05< 0.05, **p <0.01<0.01< 0.01, ***p <0.001<0.001< 0.001 compared to best baseline (two-tailed t-test)
TABLE XV: Performance across code synthesis and debugging tasks (mean ± std over 5 runs). Note that while Chronos shows average performance on pure code generation tasks (HumanEval, MBPP), it dramatically outperforms all models on debugging-specific metrics.
V-H Comprehensive Debugging Performance Analysis
To provide a complete picture of debugging capabilities, Table XVI presents detailed metrics across all evaluation dimensions:
TABLE XVI: Comprehensive debugging performance comparison across all key metrics
Model/System Root Cause Fix Valid Cross-File MRR Regression Debug Time to
Precision (%) Accuracy (%) Hit Rate (%) Score Avoid (%) Cycles Fix (min)
General-Purpose LLMs
GPT-4.1 31.2±2.1 13.8±1.2 42.3±3.1 0.28 67.2±4.2 8.3±2.1 47.2±12.3
Claude 4 Opus 34.7±2.3 14.2±1.3 45.8±3.3 0.31 69.8±4.0 7.9±1.9 44.6±11.8
Gemini 2.5 Pro 29.8±2.0 13.9±1.2 39.2±2.9 0.26 65.3±4.4 9.1±2.3 51.3±13.7
DeepSeek V3 27.3±1.9 12.1±1.1 37.6±2.8 0.24 63.7±4.5 9.7±2.5 54.8±14.2
Enhanced RAG Systems
GPT-4.1 + HyDE 38.9±2.5 22.3±1.8 51.2±3.5 0.36 74.2±3.8 6.2±1.5 38.7±9.4
Claude 4 + Self-RAG 41.2±2.6 24.7±1.9 53.7±3.6 0.39 76.1±3.6 5.8±1.4 36.2±8.9
Gemini + GraphRAG 43.8±2.7 28.9±2.1 58.3±3.7 0.42 78.3±3.4 5.3±1.3 33.4±8.1
Specialized Debugging Tools
SWE-agent 47.3±2.8 22.9±1.7 61.2±3.8 0.44 81.2±3.2 4.7±1.1 31.2±7.6
AutoCodeRover 52.1±2.9 31.2±2.2 64.8±3.9 0.48 83.7±3.0 4.2±1.0 28.3±6.9
LangGraph + ReAct 44.6±2.7 31.2±2.2 56.9±3.7 0.41 79.8±3.3 5.1±1.2 32.6±7.9
Chronos 87.4±1.2 65.3±1.4 89.2±1.3 0.82 94.6±0.9 2.2±0.4 14.7±3.2
All metrics averaged over 5,000 debugging scenarios. ± indicates standard error. MRR Score: Multi Random Retrieval benchmark (0-1 scale).
Key insights from comprehensive analysis:
  • Root Cause Precision: Chronos achieves 87.4% accuracy in identifying the true source of bugs, 2.5x better than the best baseline
  • Cross-File Retrieval: 89.2% hit rate demonstrates AGR’s effectiveness at traversing dependencies
  • Regression Avoidance: 94.6% of Chronos fixes don’t introduce new bugs, critical for production deployment
  • Efficiency: Average 2.2 debug cycles and 14.7 minutes to fix, showing rapid convergence
V-I Comparison with Agentic Code Tools
@@
TABLE XVII: Comparison of Chronos with modern agentic code tools (2025), sorted by debugging success rate.
Category Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
IDE-Integrated Cursor IDE + 32K Session Agent Mode Yes No 4.2%
Windsurf Cascade Tech Session Write Mode Yes No 5.1%
Code Assistants GitHub Copilot X 16K tokens None No Limited No 5.3%
Warp.dev Terminal Session No Limited Yes 7.3%*
CLI Tools Claude Code CLI 200K tokens Session No Yes No 6.8%
Gemini CLI 1.5M tokens None No Yes Limited 9.7%
Debugging-First Chronos Unlimited** Persistent Yes Yes Yes 65.3%
*Warp achieves 71% on SWE-bench code generation with specialized setup but only 7.3% on real-world debugging tasks (MRR benchmark). **Via AGR intelligent retrieval.
@@
V-I1 Modern Development Environment Analysis
The 2025 landscape of AI-powered development tools showcases significant innovation, yet reveals fundamental limitations in debugging capabilities:
Cursor introduced agent mode for multi-file generation but remains optimized for code completion speed over debugging accuracy. Its 4.2% debug success rate reflects prioritization of real-time suggestions over deep reasoning.
Windsurf by Codeium features Cascade technology for mapping codebases ”like a neural net,” enabling impressive multi-file edits. However, its lack of persistent memory and debugging-specific training limits effectiveness on complex bugs.
Claude Code CLI and Gemini CLI represent the evolution of terminal-based AI assistants. Claude Code CLI can use either Opus 4 or Sonnet 4 models, offering flexibility between performance and cost. While Claude Code excels at multi-file reasoning and Gemini CLI offers 1M token context, neither implements the iterative debugging loops necessary for autonomous bug resolution.
Warp.dev demonstrates interesting potential, achieving 71% on SWE-bench with specialized configuration, but this performance doesn’t generalize to real-world debugging scenarios where it achieves only 7.3% success.
Agentic and Reasoning Models: Despite recent advances, even sophisticated approaches fail at debugging. Claude 4 Opus’s agent mode with chain-of-thought reasoning achieves only 16.8% debug success, marginally better than its baseline. Reasoning-specific models like o1-preview (18.2%) and o1-mini (15.7%) show similar limitations. Web search integration provides modest improvements: Claude with web search reaches 19.3%, while GPT-4.1 with browsing achieves 17.6%. These results confirm that debugging requires more than enhanced reasoning or information access, it demands specialized architecture, persistent memory, and iterative refinement capabilities that current models lack.
V-I2 Comparative Debugging Loop Analysis
We analyze the debugging approaches of various systems to understand why Chronos achieves superior performance:
TABLE XVIII: Debugging loop characteristics comparison across systems
System Loop Type Avg Iter Test Exec Memory Backtrack Success
Chronos Autonomous 7.8 Persistent 65.3%
Claude 4 Opus Single-shot 1.2 $\times$ Session $\times$ 14.2%
GPT-4.1 Limited retry 2.1 $\times$ Session $\times$ 13.8%
Cursor User-driven 3.5 Manual Session $\times$ 4.2%
Windsurf Multi-file 2.8 Manual Session $\times$ 6.1%
Amazon Q Guided 4.2 Session Limited 49.0%
GitHub Copilot Suggestion 1.0 $\times$ None $\times$ 8.7%
Warp.dev Config-based 3.1 Session $\times$ 7.3%
Key observations:
  • Iteration Depth: Chronos averages 7.8 iterations vs 1-4 for others, enabling deeper exploration
  • Test Integration: Only Chronos, Amazon Q, and Warp.dev execute tests automatically
  • Backtracking: Chronos uniquely supports full hypothesis backtracking when fixes fail
  • Memory Persistence: All competitors use session-based memory, losing context between runs
V-J Quantitative Analysis: Debug Cycles and Convergence Rates
@@
GPT-4.1Claude 4 OpusDeepSeek V3Chronos02224446668884.84.84.84.54.54.54.24.24.22.22.22.2ModelAvg. Cycles to Fix
@@
Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles, underscoring its benefit for continuous, automated codebase reliability.
V-K Real-World Debugging Scenarios: Detailed Case Analysis
@@
V-K1 Case Study 1: Cross-Module Null Pointer Exception
@@
Technical Analysis:
  • Initial Context: 3,247 tokens from stack trace, 2 affected files
  • AGR Retrieval: k=1 found 3 auth commits, k=2 expanded to 7 related files, k=3 included test cases
  • Confidence Progression: 32% $\rightarrow$ 67% $\rightarrow$ 91% (terminated at k=3)
@@
    Semantic Search (142ms): Retrieved commits with ”authentication” keyword via Memory Engine
@@
    Data Flow Analysis (287ms): Traced AuthToken propagation through 5 call sites
@@
    Pattern Recognition (95ms): Identified missing null-safety pattern post-refactor
@@
    Impact Analysis (178ms): Located 2 similar vulnerabilities in UserService and AdminService
@@
Null Pointer Fix in Authentication Flow
// In AuthService.java
public AuthToken refreshToken(String userId) {
AuthToken token = tokenCache.get(userId);
if (token == null || token.isExpired()) {
token = authProvider.generateToken(userId);
+ if (token != null) {
+ tokenCache.put(userId, token);
return token;
// In ExportService.java
public void exportUserData(String userId) {
AuthToken token = authService.refreshToken(userId);
+ if (token == null) {
+ throw new AuthenticationException(
+ "Failed to authenticate user for export");
// ... rest of export logic
Validation Results:
  • Test Coverage: 47/47 existing tests passed, 3 new edge case tests generated
  • Performance Impact: No measurable latency increase (<1<1< 1ms overhead)
  • Fix Iterations: Converged in 2 cycles (initial fix $\rightarrow$ test generation $\rightarrow$ validation)
  • Total Time: 1.7 seconds from bug report to validated fix
V-K2 Case Study 2: Async Race Condition in Message Queue
@@
Technical Analysis:
  • Initial Context: 8,432 tokens (logs, metrics, queue implementation)
  • AGR Retrieval: k=4 required due to distributed system complexity
  • Pattern Matching: Found similar fix in commit a3f8b2c via Memory Engine
Chronos Debugging Process:
  1. 1.
    Log Analysis (523ms): Correlated message IDs with acknowledgment timestamps
  2. 2.
    Concurrency Detection (1,247ms): Identified non-atomic operation between ack() and releaseConnection()
  3. 3.
    Historical Search (432ms): Retrieved 3 similar race conditions from Memory Engine
  4. 4.
    Solution Synthesis (298ms): Adapted previous synchronization pattern to current architecture
@@
Race Condition Fix in Message Queue
// MessageProcessor.java
private void processMessage(Message msg) {
try {
handler.process(msg);
- connectionPool.returnConnection(conn);
- msg.acknowledge();
+ // Fix: Acknowledge before returning connection
+ msg.acknowledge();
+ connectionPool.returnConnection(conn);
} catch (Exception e) {
+ // Ensure connection returned even on error
+ connectionPool.returnConnection(conn);
msg.nack();
throw e;
@@
V-K3 Complete Debugging Trace: Memory Persistence Example
To demonstrate Chronos’s full debugging workflow and memory retention, we present a complete trace from our evaluation:
[…]
PDM Match: 87% similarity to ”Redis timestamp precision”
Chronos: Retrieved previous fix, adapted to session context
Time to Fix: 47 seconds (vs 3.2 minutes for original)
This trace demonstrates how PDM enables Chronos to learn from past debugging sessions, dramatically reducing fix time for similar issues.
V-L Performance at Scale: Million-Token Context Evaluation
@@
GPT-4.1 128K tokens 15.7±2.2% 7.8±1.6% 13.2±1.9% 12.2±1.8%
Claude 4 Opus 200K tokens 16.2±2.1% 8.3±1.5% 14.1±1.9% 12.9±1.8%
DeepSeek V3 128K tokens 13.5±2.2% 6.9±1.6% 11.8±2.0% 10.7±1.9%
Gemini 2.5 Pro 1M tokens 17.1±2.1% 9.2±1.5% 15.3±1.8% 13.9±1.7%
Chronos Unlimited* 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
*Via AGR. ***p <0.001<0.001< 0.001 compared to Gemini 2.5 Pro (paired t-test, n=100 tasks per category)
TABLE XIX: Performance on debugging tasks requiring extensive context. Despite models having up to 1M tokens, intelligent retrieval and debug-specific training enable Chronos to achieve 5x better performance.
The results demonstrate that raw context size alone cannot solve debugging. Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 5x.
V-M Detailed Performance Analysis
@@
GPT-4.1 88.7% 17.3% 5.8% 8.2% 24.6% 11.3%
Claude 4 Opus 91.2% 18.9% 6.3% 9.1% 26.8% 12.7%
DeepSeek V3 87.9% 16.2% 4.9% 7.6% 23.1% 10.5%
Qwen2.5-32B 86.8% 15.7% 4.5% 7.2% 22.3% 9.8%
Gemini 2.5 Pro 89.3% 17.8% 5.7% 8.7% 25.3% 11.9%
@@
TABLE XX: Success rates by bug category. While frontier models show incremental improvements over their predecessors, Chronos demonstrates 3-10x better performance on complex bug types through debug-specific training.
[…]
GPT-4.1 18.5% 13.2% 7.8% 3.1%
Claude 4 Opus 21.7% 15.8% 9.2% 4.3%
DeepSeek V3 19.2% 13.9% 8.1% 3.5%
Gemini 2.5 Pro 24.1% 17.3% 11.5% 5.2%
@@
TABLE XXI: Debugging success rates by repository size, demonstrating Chronos’s scalability.
V-N Multi-Code Association Retrieval Performance
@@
TABLE XXII: Multi-code association retrieval performance across different debugging contexts.
V-O Efficiency Metrics: Cost, Latency, and Resource Usage
@@
Metric GPT-4.1 Claude 4 Opus Gemini 2.5 Pro Chronos Human Dev
Avg. Time to Fix 68.5s 64.2s 59.8s 134.7s 2.4 hours
@@
Cost per Bug \$0.52 \$0.58 \$0.72 \$0.89 \$180
Success Rate 13.8% 14.2% 13.9% 65.3% 94.2%
Effective Cost* \$3.77 \$4.08 \$5.18 \$1.36 \$191
@@
TABLE XXIII: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
V-O1 Inference Time Breakdown
@@
V-O2 Return on Investment Analysis
@@
V-O3 Detailed Inference Cost and Latency Analysis
We conducted comprehensive benchmarking of inference costs and latency across different bug complexity levels:
TABLE XXIV: Inference cost and latency breakdown by bug complexity and system
System Simple Bugs Medium Bugs Complex Bugs
Cost (\$) Latency (s) Cost (\$) Latency (s) Cost (\$) Latency (s)
GPT-4.1 (128K) 0.18 12.3 0.52 68.5 1.24 187.2
Claude 4 Opus 0.21 14.7 0.58 64.2 1.41 201.3
Gemini 2.0 Pro 0.31 8.9 0.72 59.8 1.93 156.4
Amazon Q Dev 0.42 21.3 0.91 94.7 2.14 234.6
Chronos (Ours) 0.34 31.2 0.89 134.7 1.78 298.4
- Retrieval only 0.08 7.8 0.19 23.4 0.41 52.3
- LLM inference 0.21 18.9 0.56 67.8 1.12 156.8
- Test execution 0.03 3.1 0.09 31.2 0.18 67.2
- Memory update 0.02 1.4 0.05 12.3 0.07 22.1
Cost-Effectiveness Analysis (Cost / Success Rate)
GPT-4.1 1.30 - 3.77 - 8.99 -
Claude 4 Opus 1.48 - 4.08 - 9.93 -
Gemini 2.0 Pro 2.23 - 5.18 - 13.88 -
Amazon Q Dev 0.86 - 1.86 - 4.37 -
Chronos 0.52 - 1.36 - 2.73 -
Key observations:
  • Latency vs Accuracy Trade-off: Chronos’s higher latency (2-3x) is offset by 4-5x better success rates
  • Cost Scaling: Complex bugs show sublinear cost increase (5.2x) despite exponential difficulty
  • Component Distribution: LLM inference dominates cost (63%) while retrieval dominates initial latency
  • Parallelization Opportunity: Test execution (23% of latency) can be parallelized for 1.3x speedup
VI ANALYSIS: WHY CHRONOS SUCCEEDS WHERE OTHERS FAIL
VI-A The Debugging Specialization Hypothesis: Evidence and Implications
The stark performance gap between Chronos (65.3% debug success) and state-of-the-art general models ($\leq$15%) reveals a fundamental truth: debugging is not simply an extension of code generation. While models like Claude 4 Opus (72.5% on SWE-bench) and GPT-4.1 (54.6%) excel at writing new code, debugging demands distinct capabilities:
  1. 1.
    Temporal Reasoning: Understanding how code evolved through commits, why certain patterns were introduced, and which changes correlate with bug emergence. General models lack this historical perspective.
  2. 2.
    Multi-Modal Signal Integration: Debugging requires synthesizing error traces, logs, test failures, and code, a fundamentally different task than generating syntactically correct code from specifications.
  3. 3.
[…]
  4. 4.
[…]
VI-B Key Architectural Insights: Memory, Iteration, and Context